OpenStudy (anonymous):

What is the solution set for -x^2+9>?

OpenStudy (anonymous):

\[-x^2+9>\]

OpenStudy (anonymous):

Yea but do you know what the solution set is?

OpenStudy (anonymous):

> than what? ...

OpenStudy (anonymous):

Oops. Sorry Kreshnik. >0

OpenStudy (lgbasallote):

go @Kreshnik ! hehe

OpenStudy (anonymous):

that's better ;) .... we have: \[\Large -x^2+9>0\] \[\Large 9-x^2>0\quad \longrightarrow \[3^2-x^2>0\]\] now we use this formula... \[\Large a^2-b^2=(a-b)(a+b)\] and we have.. \[\Large 3^2-x^2>0 \longrightarrow (3-x) (3+x )>0\] from here we have two parts 3-x>0 or 3+x>0 so we get that... -x>-3 or x>-3 multiply left inequality by -1 and swap the sight > into < (everytime you multiply\divide by a negative number in inequalities you change sign !) and finally we have: x<3 or x>-3 so \[\LARGE x\in (-\infty ,3) \cup (-3, +\infty)\] from here we get that... \[\LARGE x\in (-3,-2,-1,0,1,2,3)\]

OpenStudy (anonymous):

Wow you're great Kreshnik. I really appreciate all of your help.

OpenStudy (anonymous):

@lgbasallote :P am I right or not ? :P

OpenStudy (anonymous):

@mathsux4real Glad to help ;) (I got to go to school, I'm out ! see you later ;) )

OpenStudy (anonymous):

Cool, see ya around!

OpenStudy (lgbasallote):

yes :D very good job @Kreshnik *thumbs up* ;D

OpenStudy (anonymous):

bye ! ;) @lgbasallote thanks ;)

OpenStudy (anonymous):

actually it should be... \[\LARGE x\in (-\infty,3)\cap(-3,+\infty)\] the rest is ok.